# PhysGM: Large Physical Gaussian Model for Feed-Forward 4D Synthesis

### v1

@@ Introduction
We focus on three key physical attributes that govern the behavior of deformable objects: a material class $C$, Young’s modulus $E$ (stiffness), Poisson’s ratio $\nu$ (compressibility) and density $\rho$. We categorize materials into $N_{c}$ discrete classes, where each class corresponds to a specific constitutive model used in the subsequent MPM simulation.
From the latent feature vector $\mathbf{z}$, we predict these properties using two specialized heads: A classification head $f_{\text{material}}$ to determine the material class; A multi-layer perceptron (MLP) head, $f_{\text{phys}}$ predicting the normal distribution for the continuous physical properties $\boldsymbol{\theta}=[E,\nu,\rho]^{T}$. Specifically, $f_{\text{phys}}$ outputs the mean $\boldsymbol{\mu}_{\theta}$ and log-variance $\log\boldsymbol{\sigma}_{\theta}^{2}$ for these properties:
$(\boldsymbol{\mu}_{\theta},\log\boldsymbol{\sigma}_{\theta}^{2})=f_{\text{phys}}(\mathbf{z}).$ (2)
@@
Figure 3: Preference calculation. We use SAM-2 for instance segmentation and CoTracker-3 for trajectory extraction across the GT and simulated videos. The extracted point tracks quantify the fidelity of each candidate to the GT, yielding a ranked preference tuple.
Probabilistic Appearance Modeling
For each of the $N$ Gaussians comprising the scene, we predict a full set of its defining parameters, denoted as $\boldsymbol{\psi}$. Similar to Section Probabilistic Modeling, a dedicated network head $f_{\text{gauss}}$ processes the latent feature vector $\mathbf{z}$ to output the parameters of a Gaussian distribution for the scene’s attributes:
$(\boldsymbol{\mu}_{\psi},\log\boldsymbol{\sigma}_{\psi}^{2})=f_{\text{gauss}}(\mathbf{z}).$ (4)
This defines the conditional probability of the full set of Gaussian parameters given the initial image:
$P(\boldsymbol{\psi}|I)=\mathcal{N}(\boldsymbol{\psi}|\boldsymbol{\mu}_{\psi},\text{diag}(\boldsymbol{\sigma}_{\psi}^{2})).$ (5)
During pre-training, the network is optimized via a dual objective. It jointly minimizes the Negative Log-Likelihood (NLL) of the ground-truth physical and geometric parameters under the model’s predicted distributions, and a photometric reconstruction loss between the image rendered from the generated Gaussians and the ground-truth view. This joint supervision ensures that the learned generative prior is not only physically accurate but also capable of producing high-quality scene representations.
At inference time, we sample from these learned distributions to obtain a specific instance of the scene’s parameters: $\boldsymbol{\psi}_{\text{sampled}}\sim P(\boldsymbol{\psi}|\mathbf{z}),\boldsymbol{\theta}_{\text{sampled}}\sim P(\boldsymbol{\theta}|\mathbf{z})$. This probabilistic formulation is crucial, as it allows the model to generate diverse Gaussian and physical parameters, enabling the subsequent preference-based refinement with GT videos.
@@
Physics-based Dynamics via MPM
To simulate physically based dynamics, we employ the Material Point Method (MPM) (Jiang et al. 2016; Stomakhin et al. 2013), a hybrid Lagrangian-Eulerian approach that discretizes an object into a set of material points. Each point $p$ carries its own physical state, including mass $m_{p}$, position $\mathbf{x}_{p}$, velocity $\mathbf{v}_{p}$, and the affine velocity matrix $\mathbf{C}_{p}$. Crucially, it also tracks the deformation gradient $\mathbf{F}_{p}$, which maps vectors from the material’s rest configuration to its current deformed state.
@@
Following PhysGaussian (Xie et al. 2024), we directly couple this physical simulation with our 3D Gaussian representation. We establish a one-to-one correspondence between each material point and a Gaussian primitive. The updated particle position $\mathbf{x}_{p}$ directly defines the Gaussian’s mean $\boldsymbol{\mu}$. More importantly, the deformation gradient $\mathbf{F}_{p}$ dictates the Gaussian’s anisotropic shape and orientation. We perform a polar decomposition on $\mathbf{F}_{p}$ to factor it into a rotation matrix $\mathbf{R}_{p}$ and a symmetric positive semi-definite stretch tensor $\mathbf{S}_{p}$, $\mathbf{F}_{p}=\mathbf{R}_{p}\mathbf{S}_{p}$. The resulting rotation $\mathbf{R}_{p}$ and the diagonal elements of the stretch tensor $\mathbf{S}_{p}$ are then used to define the Gaussian’s rotation matrix $\mathbf{R}$ and scaling vector $\mathbf{s}$, respectively. These are then used to construct the covariance matrix $\boldsymbol{\Sigma}$, thus ensuring the rendered geometry precisely mirrors the physically simulated deformation.
Figure 5: Ablation results of DPO: The results indicate that after the two-stage DPO training, the model predicts physical attributes with greater accuracy, enabling the generation of 4D videos that exhibit higher physical fidelity.
Preference-based Fine-tuning with DPO
While supervised pre-training provides a robust physical prior, it may not fully capture the subtle dynamics that lead to high perceptual quality. To bridge this gap, we introduce a fine-tuning stage using Direct Preference Optimization (DPO) (Rafailov et al. 2023). This approach allows us to refine our generative model using feedback from the full physics simulation and rendering pipeline without requiring this pipeline to be differentiable.
We treat the pre-trained model as a fixed reference policy, $\pi_{\text{ref}}$. The model being optimized, $\pi_{\omega}$, is then refined using a dataset of preference pairs $\mathcal{D}=\{(\mathbf{z},\boldsymbol{\phi}_{w},\boldsymbol{\phi}_{l})\}$. For each scene context $\mathbf{z}$, we perform the following steps to create a preference pair. First, we draw a set of $K$ candidate parameter vectors $\{\boldsymbol{\phi}_{1},\dots,\boldsymbol{\phi}_{K}\}$ from our current policy, where $\boldsymbol{\phi}_{k}\sim\pi_{\omega}(\cdot|\mathbf{z})$. Each $\boldsymbol{\phi}$ consists of the predicted physical properties $\theta$ and appearance parameters $\psi$. For each candidate $\boldsymbol{\phi}_{k}$, we run the MPM simulation and render the resulting 3D Gaussian sequence to produce a short video clip $V_{k}$. We compare each generated clip $V_{k}$ to the ground-truth video $V_{gt}$ using a perceptual distance metric $d(V_{k},V_{gt})$. The parameter set $\boldsymbol{\phi}_{k}$ that yields the lowest perceptual distance is designated the “winner” $\boldsymbol{\phi}_{w}$. Another sample with a higher distance is chosen as the “loser” $\boldsymbol{\phi}_{l}$.
@@
$L_{\text{DPO}}(\pi_{\omega},\pi_{\text{ref}})=-\mathbb{E}_{(\mathbf{z},p_{w},p_{l})\sim D}[\log\sigma(p_{1}-p_{2})],$ (10)
$p_{1}=\beta\log\frac{\pi_{\omega}(p_{w}|\mathbf{z})}{\pi_{\text{ref}}(p_{w}|\mathbf{z})},\ p_{2}=\beta\log\frac{\pi_{\omega}(p_{l}|\mathbf{z})}{\pi_{\text{ref}}(p_{l}|\mathbf{z})},$ (11)
where $\beta$ is a temperature parameter controlling the optimization strength. By optimizing this objective, we effectively steer the learned distribution towards regions of the parameter space that produce perceptually and physically superior simulations, without the need for optimization upon complex differentiable MPM and rendering process.
@@
OmniPhysGS 0.2149 86% 0.2291 78% 0.2135 74% 0.1834 80% 0.2047 62% 0.2091 76%
DreamPhysics 0.2273 88% 0.2459 64% 0.2437 68% 0.2071 72% 0.2217 78% 0.2291 74%
PhysGM (w/o DPO) 0.2698 58% 0.2700 54% 0.2547 60% 0.2541 52% 0.2980 62% 0.2693 57%
PhysGM (w/ DPO) 0.2732 - 0.2774 - 0.2691 - 0.2548 - 0.2997 - 0.2748 -
Table 1: Quantitative comparisons. We evaluate our method and baseline models on 5 different material types. Evaluation is based on the $\text{CLIP}_{sim}$ score (higher is better $\uparrow$) and UPR (lower is better $\downarrow$).
PhysAssets Dataset
We introduce PhysAssets, a multi-modal dataset of over 24,000 3D objects designed to establish a robust mapping between object geometry and physical properties. The 3D assets were aggregated from established datasets, including Objaverse (Deitke et al. 2023), OmniObject3D (Wu et al. 2023), ABO (Collins et al. 2022), and GSO (Downs et al. 2022). To annotate physical attributes, we developed a pipeline leveraging a Vision-Language Model (VLM) SigLIP-2 (Tschannen et al. 2025). For each object, we compute the cross-modal similarity between multi-view images and predefined textual descriptors for material properties. The ground-truth (GT) video is generated for each object using the FramePack model (Zhang and Agrawala 2025) with text instructions on the detailed description of the physical properties of the object. We automate preference labeling $d(V_{k},V_{gt})$ via a comparative analysis pipeline, as shown in Figure 3. This involves using SAM-2 (Kirillov et al. 2023; Ravi et al. 2024) for instance segmentation and CoTracker-3 (Karaev et al. 2024b, a) for trajectory extraction across the GT and simulated videos. The extracted point tracks then quantify the fidelity of each candidate relative to the GT, yielding a ranked preference tuple. More details on video generation and preference calculation process can be found in supplementary materials.
Experiments
In this section, we conduct a comprehensive comparison against several baseline methods, evaluating the stability of fine-grained physical properties. Furthermore, we perform ablation studies on our two-stage training strategy to validate its effectiveness.
Experimental settings
Dataset
We train and evaluate our model on our newly created PhysAssets Dataset. For qualitative visualizations, we show results on the test set from our dataset as well as on in-the-wild images to demonstrate generalization.
Baselines
We compare our method against two baseline models in physics-based dynamic generation: OmniPhysGS (Lin et al. 2025) and DreamerPhysics (Huang et al. 2025). These methods represent the dominant paradigm of per-scene optimization, where physical properties are learned by distilling knowledge from video models using Score Distillation Sampling (SDS). They provide strong benchmarks for simulation quality, albeit at a significant computational cost.
@@
The physical properties predicted by our model directly drive the subsequent Material Point Method (MPM) simulation. The predicted material class determines which constitutive model is employed to govern the object’s dynamic behavior. For instance, a prediction of ‘rubber’ would select a Neo-Hookean constitutive model. The predicted Young’s modulus ($E$) and Poisson’s ratio ($\nu$) then serve as the specific parameters for this chosen model. We use a fixed sub-step time of $2\times 10^{-5}$s and a frame time of $4\times 10^{-2}$s, generating 50 frames per sequence. Additional simulation parameters, such as boundary conditions, are detailed in the Appendix.
@@
We initialize our model with the weight of LGM (Tang et al. 2024). We train our model for 12 hours on 8 NVIDIA A800 GPUs, with a batch size of 8 per GPU. The training process is conducted in two stages. In the pre-training phase, we jointly optimize our model for physics and Gaussian prediction under the supervision of PhysAssets. In the fine-tuning stage, we apply DPO to guide the model to generate more physically realistic dynamics. During this stage, we freeze the backbone of the model, fine-tuning only the predictive heads for Gaussian parameters and physics properties.
@@
We conduct a comprehensive evaluation using two complementary metrics: an objective, model-based score and a subjective, human-based assessment. (1) $\text{CLIP}_{sim}$ (Radford et al. 2021) quantifies the semantic similarity between the rendered visual outcomes and their corresponding textual descriptions of the physical phenomena. (2) User Preference Rate (UPR). We conduct a Two-Alternative Forced-Choice (2AFC) user study. The participants are presented with side-by-side results from our method and a competitor, asked to select the more realistic or visually appealing one. UPR is defined as the percentage of trials where our method was chosen.
Results and Analysis
@@
We begin by showcasing the core capability of optimization-free 4D generation. As illustrated in Figure 4, our model takes a single input image and, in a single feed-forward pass, generates a complete 3D Gaussian representation along with its corresponding physical properties. This allows for the production of a full, physically based simulation in under one minute.
For a quantitative comparison, we evaluate PhysGM against two optimization-based baselines, OmniPhysGS and DreamerPhysics. Our method outperforms both baselines on both metrics across a diverse range of materials. This result demonstrates that our feed-forward approach does not merely trade quality for speed. Instead, by learning a robust physical prior from a large dataset, PhysGM can surpass the perceptual realism and physical plausibility of slower, per-scene optimization techniques.
@@
To validate the effectiveness of our proposed two-stage training strategy, we conduct an ablation study focused on the DPO fine-tuning stage. We compare our full PhysGM model against a baseline variant, “w/o DPO,” which is trained exclusively with supervised pre-training and does not undergo DPO refinement.
The results presented in Table 1 and visualized in Figure 5 clearly demonstrate the critical role of the DPO stage. The full model achieves substantially higher scores across all categories in both metrics. The qualitative comparison is particularly revealing. While the pretraining-only model might predict a plausible but overly rigid material, the DPO-refined model learns to predict more accurate physical parameters. This leads to simulations with significantly more natural deformations. These findings confirm that our DPO refinement stage is essential for transforming a statistically sound prior into a perceptually better generative model. It effectively uses the feedback from the entire simulation pipeline, steering the model towards outputs that are not just statistically likely but are physically and visually more compelling.
Conclusion

### v2

@@ 3.2 Model Architecture
We predict three physical attributes that determine kinetic behavior of the particles: a material class $C$, Young’s modulus $E$ (stiffness) and Poisson’s ratio $\nu$ (compressibility). We categorize materials into $N_{c}$ classes, where each class corresponds to a specific constitutive model used in the subsequent MPM simulation. From the three global tokens $\mathbf{g}_{k}$, we predict these properties using two specialized heads: A classification head $f_{\text{material}}$ to determine the material class; two regression heads $f_{\text{phys}}$ predicting the distribution for the continuous physical properties $E,\nu$ respectively. Specifically, $f_{\text{phys}}$ outputs the mean $\boldsymbol{\mu}_{\theta}$ and log-variance $\log\boldsymbol{\sigma}_{\theta}^{2}$ for these properties:
$(\boldsymbol{\mu}_{\theta},\log\boldsymbol{\sigma}_{\theta}^{2})=f_{\text{phys}}(\mathbf{g}_{k}).$ (3)
@@
At inference time, we sample from these learned distributions to obtain the scene’s parameters: $\boldsymbol{\theta}_{\text{sampled}}\sim P(\boldsymbol{\theta}|I)$. This probabilistic formulation is crucial, as it allows the model to generate diverse physical parameters, enabling the subsequent preference-based refinement with GT videos.
Figure 3: Preference calculation. We use SAM-2 [ravi2024sam] for segmentation and CoTracker-3 for trajectory extraction across the GT and simulated videos. The extracted point tracks quantify the fidelity of each candidate to the GT, yielding a ranked preference tuple.
@@
3.3 Physics-based Dynamics via MPM
To simulate physics-based dynamics, we employ the Material Point Method (MPM) [jiang2016material, stomakhin2013material], a hybrid Lagrangian-Eulerian approach that discretizes an object into a set of material points. Each point $p$ carries its own physical state, including mass $m_{p}$, position $\mathbf{x}_{p}$, velocity $\mathbf{v}_{p}$, and the affine velocity matrix $\mathbf{C}_{p}$. Crucially, it also tracks the deformation gradient $\mathbf{F}_{p}$, which maps vectors from the material’s rest configuration to its current deformed state.
@@ 3.3 Physics-based Dynamics via MPM
Following PhysGaussian [xie2024physgaussian], we directly couple this physical simulation with our 3D Gaussian representation. The MPM simulation is driven by the physical properties predicted by our model. We establish a one-to-one correspondence between each material point and a Gaussian primitive. The updated particle position $\mathbf{x}_{p}$ directly defines the Gaussian’s mean $\boldsymbol{\mu}$. More importantly, the deformation gradient $\mathbf{F}_{p}$ dictates the Gaussian’s anisotropic shape and orientation. We perform a polar decomposition on $\mathbf{F}_{p}$ to factor it into a rotation matrix $\mathbf{R}_{p}$ and a symmetric positive semi-definite stretch tensor $\mathbf{S}_{p}$, $\mathbf{F}_{p}=\mathbf{R}_{p}\mathbf{S}_{p}$. The resulting rotation $\mathbf{R}_{p}$ and the diagonal elements of the stretch tensor $\mathbf{S}_{p}$ are then used to define the Gaussian’s rotation matrix $\mathbf{R_{m}at}$ and scaling vector $\mathbf{s}$, respectively. These are used to construct the covariance matrix $\boldsymbol{\Sigma}$, ensuring the rendered geometry mirrors the physically simulated deformation.
3.4 Preference-based Fine-tuning with DPO
While supervised pre-training provides a robust physical prior, it may not fully capture the subtle dynamics that lead to high perceptual quality. To bridge this gap, we introduce a fine-tuning stage using Direct Preference Optimization (DPO) [rafailov2023direct]. This approach enables refining our generative model using feedback from physics simulation and rendering pipeline without requiring differentiability.
We treat the pre-trained model as a fixed reference policy, $\pi_{\text{ref}}$. The model being optimized, $\pi_{\omega}$, is then refined using a dataset of preference pairs $\mathcal{D}=\{(\mathbf{z},\boldsymbol{\phi}_{w},\boldsymbol{\phi}_{l})\}$. For each scene context $\mathbf{z}$, we perform the following steps to create a preference pair. First, we draw a set of $K$ candidate parameter vectors $\{\boldsymbol{\phi}_{1},\dots,\boldsymbol{\phi}_{K}\}$ from our current policy, where $\boldsymbol{\phi}_{k}\sim\pi_{\omega}(\cdot|\mathbf{z})$. Each $\boldsymbol{\phi}$ consists of the predicted physical properties $\theta$ and fixed appearance parameters $\psi$. For each candidate $\boldsymbol{\phi}_{k}$, we run the MPM simulation and render the resulting 3D Gaussian sequence to produce a short video clip $V_{k}$. We compare each generated clip $V_{k}$ to the ground-truth video $V_{gt}$ using a perceptual distance metric $d(V_{k},V_{gt})$.The parameter set $\boldsymbol{\phi}_{k}$ that yields the lowest perceptual distance is designated the “winner” $\boldsymbol{\phi}_{w}$. Another sample with the highest distance is chosen as the “loser” $\boldsymbol{\phi}_{l}$.
@@ 3.4 Preference-based Fine-tuning with DPO
$L_{\text{DPO}}(\pi_{\omega},\pi_{\text{ref}})=-\mathbb{E}_{(\mathbf{z},\boldsymbol{\phi}_{w},\boldsymbol{\phi}_{l})\sim D}[\log\sigma(p_{1}-p_{2})],$ (9)
$p_{1}=\beta\log\frac{\pi_{\omega}(\boldsymbol{\phi}_{w}|\mathbf{z})}{\pi_{\text{ref}}(\boldsymbol{\phi}_{w}|\mathbf{z})},\ p_{2}=\beta\log\frac{\pi_{\omega}(\boldsymbol{\phi}_{l}|\mathbf{z})}{\pi_{\text{ref}}(\boldsymbol{\phi}_{l}|\mathbf{z})},$ (10)
where $\beta$ is a temperature parameter controlling the optimization strength. By optimizing this objective, we steer the learned distribution towards regions of the parameter space that produce perceptually and physically superior simulations, without the need for optimization upon complex differentiable MPM and rendering process.
3.5 PhysAssets Dataset
We introduce PhysAssets, a large dataset of over 50,000 3D objects aggregated from established datasets, including Objaverse [deitke2023objaverse], OmniObject3D [wu2023omniobject3d], ABO [collins2022abo] and HSSD [khanna2023hssd]. To annotate the physical attributes, we developed a pipeline leveraging a large Multimodal Large Language Model (MLLM) Qwen3VL [qwen3technicalreport]. For each object, we first obtain its corresponding material category directly using a well-designed prompt with multi-view images. Concurrently, we extract the physical property parameters of the object through predefined textual descriptions of Young’s modulus and Poisson’s ratio. For each object, we generate ground-truth (GT) videos using Framepack [zhang2025packing], a physics-based simulation framework. The simulation is conditioned on text instructions that specify the detailed physical properties of the object, ensuring realistic dynamic behaviors.
To enable DPO training, we automate preference labeling $d(V_{k},V_{gt})$ via a comparative analysis pipeline, as shown in Figure 3. This involves using SAM-2 [kirillov2023segment, ravi2024sam] for instance segmentation and CoTracker-3 [karaev2024cotracker, karaev24cotracker3] for trajectory extraction across the GT and simulated videos. The extracted point tracks then quantify the fidelity of each candidate relative to the GT, yielding a ranked preference tuple. Comprehensive details on the composition of the data set, video generation pipeline, and the preference calculation process are provided in the supplementary materials.
4 Experiments
In this section, we conduct a comprehensive comparison against several baseline methods, evaluating the effectiveness and efficiency of our model. Furthermore, we perform ablation studies on our two-stage training strategy.
4.1 Experimental settings
Dataset
We train our model on our newly created PhysAssets Dataset. For qualitative visualizations, we show results on the test set from our dataset as well as on in-the-wild images to demonstrate generalization.
Baselines
We compare our method against two baseline models in physics-based dynamic generation: OmniPhysGS [lin2025omniphysgs] and DreamerPhysics [huang2025dreamphysics]. These methods represent the paradigm of per-scene optimization, where physical properties are learned by distilling knowledge from video models using Score Distillation Sampling (SDS). These methods provide strong benchmarks for simulation quality, albeit at a significant computational cost.
Table 1: Quantitative comparisons. We evaluate our method and baseline models on 5 different material types. Evaluation is based on the $\text{CLIP}_{sim}$ score (higher is better $\uparrow$) and UPR (higher is better $\uparrow$).
@@ 4.1 Experimental settings
OmniPhysGS [lin2025omniphysgs] 0.2149 5% 0.2291 12% 0.2135 8% 0.1834 9% 0.2047 16% 0.2091 10%
DreamPhysics [huang2025dreamphysics] 0.2273 16% 0.2459 11% 0.2437 23% 0.2071 18% 0.2217 18% 0.2291 17.2%
PhysGM (w/o DPO) 0.2698 30% 0.2700 33% 0.2547 31% 0.2541 26% 0.2980 30% 0.2693 30%
PhysGM (w/ DPO) 0.2732 49% 0.2774 44% 0.2691 38% 0.2548 47% 0.2997 36% 0.2748 42.8%
@@
The physical properties predicted by our model directly drive the subsequent Material Point Method (MPM) simulation. The predicted material class determines which constitutive model is employed for the object’s dynamics. For instance, a prediction of ‘rubber’ would select a Neo-Hookean constitutive model. The predicted Young’s modulus ($E$) and Poisson’s ratio ($\nu$) then serve as the specific material parameters for this chosen model. We use a fixed sub-step time of $2\times 10^{-5}$s and a frame time of $4\times 10^{-2}$s, generating 50 frames per sequence. Additional simulation parameters, such as boundary conditions, are detailed in the Appendix.
Figure 5: Other Results by PhysGM. PhysGM can demonstrate robust generalization to diverse physical interactions. It accurately simulates complex deformations like stretching and twisting, handles multi-object dynamics with varied materials, and processes real-world data, highlighting its extensibility to novel scenarios.
Figure 6: Ablation results of DPO. The results indicate that after the two-stage DPO training, the model predicts physical attributes with greater accuracy, enabling the generation of 4D videos that exhibit higher physical fidelity.
@@
Our model is trained from scratch and leverages FlashAttention v2 [dao2023flashattention2] for efficient computation, with training conducted on 32 NVIDIA A800 GPUs for 3 days and a batch size of 8 per GPU. The training process consists of two stages: in the pre-training phase, we jointly optimize the model for physics property prediction and Gaussian parameter prediction under the supervision of the PhysAssets dataset; In the fine-tuning phase, we employ Direct Preference Optimization (DPO) to further enhance the physical realism and temporal coherence of the generated dynamics, where the model’s backbone remains frozen and only the predictive heads responsible for physics properties are fine-tuned.
@@
We conduct a comprehensive evaluation using two complementary metrics: an objective, model-based score and a subjective, human-based assessment. (1) $\text{CLIP}_{sim}$ [radford2021learning] quantifies the semantic similarity between the rendered visual outcomes and their corresponding textual descriptions of the physical phenomena. (2) User Preference Rate (UPR). We conduct a four-alternative forced-choice user study; UPR is the percentage of trials in which a method is selected as the most realistic or visually appealing. More details are provided in the appendix. (3) We evaluate reconstruction quality with three metrics: PSNR measures pixel-level fidelity via the log ratio of peak signal to MSE (higher is better); SSIM [wang2004image] assesses structural similarity (0-1, higher is better); and LPIPS [zhang2018unreasonable] computes learned perceptual distance (lower is better).
4.2 Results and Analysis
@@
We demonstrate the core capability of optimization-free 4D generation in Figure 4. From a single input image, our pipeline generates a 3D Gaussian representation with physical properties via one feed-forward pass, enabling physics-based simulation in under one minute. Figure 5 present its performance on a variety of diverse and challenging scenarios in Figure 5.
For a quantitative comparison, we evaluate PhysGM against two optimization-based baselines, OmniPhysGS and DreamerPhysics. Our method outperforms both baselines on both metrics across a diverse range of materials. This demonstrates that our feed-forward approach does not trade quality for speed. By learning physical priors from large datasets, PhysGM surpasses the perceptual and physical fidelity of per-scene optimization techniques. The results presented in Table 1.
Table 2: Quantitative comparisons for multi-view synthesis on GSO dataset. We matched the baseline settings by comparing with LGM and GS-LRM, We achieve better results while using only 10% of the data compared to the GS-LRM.
Methods Res. PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
LGM [tang2024lgm] 256 21.44 0.832 0.122
Our 256 25.47 0.916 0.071
GS-LRM [zhang2024gs] 512 30.52 0.952 0.050
Ours 512 28.95 0.953 0.039
Additionally, we validate MVS effectiveness in Table 2 by comparing against Gaussian-based reconstruction methods. Qualitative results are in the appendix. Note that our model is trained only at 512 resolution.
@@
To validate the effectiveness of our two-stage training strategy, we conduct an ablation focused on the DPO fine-tuning stage. We compare the full PhysGM against a baseline variant (“w/o DPO”), which is trained only with supervised pre-train, also against prior methods.
As shown in Table 3 and Figure 6, the DPO stage is critical, the full model achieves consistently higher scores across all categories and both metrics. In effect, DPO converts a statistically sound prior into a perceptually superior generator by leveraging feedback from the full, non-differentiable simulation pipeline.
Compared with other methods, PhysGM attains better or comparable visual and physical fidelity at substantially lower optimization cost. Inference requires no per-scene optimization—only a single forward pass followed by MPM rollout—enabling end-to-end generation in under one minute, while competing approaches typically rely on iterative, scene-specific optimization. This yields a markedly better quality–time trade-off.
Table 3: Comparison with state-of-the-art methods. It can be observed that DPO achieves superior performance in generalization, inference time, and simulation quality.
Methods Training Gen. Infer time $\text{CLIP}_{sim}$
OmniPhysGS [lin2025omniphysgs] SDS $\times$ >12h 0.2091
DreamPhysics [huang2025dreamphysics] SDS $\times$ >0.5h 0.2291
Ours DPO $\checkmark$ <1min 0.2748
5 Discussion
The prohibitive computational cost of the MPM remains the primary obstacle to its use in large-scale, real-time applications, a challenge exacerbated by the lack of viable alternatives for complex physics like fluid and fracture. Critically, the persistent sim-to-real gap—stemming from inherent discrepancies between synthetic training data and physical reality, including simplified constitutive models—hinders robust real-world deployment and limits generalization capabilities. Future work must therefore prioritize two goals: developing more efficient simulation frameworks and bridging the sim-to-real gap.
6 Conclusion